# Leveraging a Cognitive Model to Measure Subjective Similarity of Human and GPT-4 Written Content

### v1

@@ 7 Case Study of IBIS: Individuals in Phishing Email Education Dataset
Semantic Similarity Qurashi et al. (2020) -1127.69 -47.69$\pm$1.19 0.46$\pm$0.11
Cosine Similarity Park et al. (2020) -2097.40 -57.26$\pm$2.27 0.52$\pm$0.10
Embedding Weighting Onan (2021) -847.56 -39.28$\pm$2.32 0.86$\pm$0.14
Embedding Pruning Manrique et al. (2023) -846.20 -37.39$\pm$2.76 0.86$\pm$0.04
Ensemble Similarity Yu et al. (2024) -812.23 -33.64$\pm$3.28 0.89$\pm$0.12
IBIS (proposed) -717.45 -30.17$\pm$3.29 0.93$\pm$0.04
@@
This individual visualization is used to highlight the variety of individual similarity metrics, as a result of individual differences in categorization accuracy, confidence, and reaction time. The KDE score of the similarities for the pruned cosine method is -36.62, and for the ensemble method it is -34.10. Note that these scores are much lower than the entire dataset scores since they are calculated using only the emails observed by the participant. Meanwhile, the IBIS metric gives a KDE score of -31.01. From this we can see that the IBIS method effectively learns the similarity measures of individual participants. These results are used for illustrative purposes, and the averages across all participants for regression accuracy, as well as the DKE score for individuals, is presented in Table 1.
An important aspect of individual similarity comparisons of the IBIS method is that it can compare emails that were not originally presented to an individual, meaning there are more embedding similarities used in the logistic regression and KDE score calculation. This comparison demonstrates the benefits of using a cognitively inspired method of modeling human participant decisions making that takes into account biases and cognitive constraints. The results is a prediction of behavior that can accurately fill in the gaps of unseen elements of the dataset that have not been observed by a participant. This method more accurately predicts the subjective similarity of participants. Importantly, this is done while initially limiting the cognitive model to observing a single decision made by these participants, and increasing this data as the participant makes more decisions.
@@ 8 Discussion
Many applications of LLMs are interested in tailoring use cases to individuals, even when little information is known about that individual. While many approaches of individualization have demonstrated success in producing outputs or representing information in an individualized manner, these have typically relied on advanced machine learning techniques. The method proposed in this work is relatively simple from a mathematical perspective, though there is a strength in its reliance on theories of cognition that underlie human learning and decision making. The result is a simple to understand and easy to implement method of calculating similarities of unseen documents using a cognitive model, which can augment datasets that contain only a small number of decisions.
The specific application we investigated is somewhat unique in that it is based on training human participants to make categorization judgements of textual information of one of two categories. However, we believe that the general method described, of augmenting subjective similarity metrics with predicted decisions from a cognitive model, could be applied onto various other scenarios. This includes settings that leverage representations formed of visual information such as $\beta$-Variational Autoencoders Higgins et al. (2016), which have been related to biological representation formation Higgins et al. (2021), and integrated into decision making systems using reinforcement learning techniques Higgins et al. (2017).
For instance, in visual learning settings VAEs have been integrated with cognitive models to predict human utility learning of abstract visual information Malloy and Sims (2024). Other integrations of Generative AI into cognitive models includes use of LLMs as a knowledge repositories within cognitive models (Kirk et al., 2023). In particular, ConceptNet (Speer et al., 2017) has previously been integrated into a cognitive model for question answering (Huet et al., 2021). Future research should investigate how additional uses of LLMs in integrations of cognitive models can aid in educational settings.
@@ 9 Limitations
The semantic similarity metric suffered from the sparsity of semantic categories in ham emails, additional annotations could raise the performance of this metric and can be explored in future work. However, this ensemble method was partially responsible for the high KDE score of the ensemble method, as it allowed for an integration of both semantic information and embedding similarity. Our IBIS method still outperformed the ensemble method suggesting that this ensemble alone does not fully address the issues of similarity methods that do not account for individual biases.
The task presented in this work of predicting whether an email is phishing or ham relies heavily on a small number of features within the email. Namely, if an email contains a link that redirects to a nefarious website, or requests personal information, then it should be labelled as phishing. While students rely on many queues to make their judgements, the annotation is in reality simple. Future work in the area of learning subjective similarity metrics should expand into domains.

### v2

@@ 7 Case Study of IBIS: Individuals in Phishing Email Education Dataset
Figure 8: Top performing similarity metrics and individual participant similarity for phishing and ham emails. Shaded region is a logistic regression. The lower value is the individual KDE score
@@
The KDE score of the similarities for the pruned cosine method is -30.76, and for the ensemble method it is -27.82. Note that these scores are much lower than the entire dataset scores since they are calculated using only the emails observed by the participant. Meanwhile, the IBIS metric gives a KDE score of -24.72. From this we can see that the IBIS method effectively learns the similarity measures of individual participants. These results are used for illustrative purposes, and the averages across all participants for regression accuracy, as well as the DKE score for individuals, is presented in Table 1.
An important aspect of individual similarity comparisons of the IBIS method is that it can compare emails that were not originally presented to an individual, meaning there are more embedding similarities used in the logistic regression and KDE score calculation. This comparison demonstrates the benefits of using a cognitively inspired method of modeling human participant decisions making that takes into account biases and cognitive constraints.
The results is a prediction of behavior that can accurately fill in the gaps of unseen elements of the dataset that have not been observed by a participant. This method more accurately predicts the subjective similarity of participants. Importantly, this is done while initially limiting the cognitive model to observing a single decision made by these participants, and increasing this data as the participant makes more decisions. This is important for the functioning of the IBL model as using too many instances in memory can slow compute performance.
@@
Semantic Similarity (Qurashi et al., 2020) -1127.69 -37.69$\pm$1.19 0.46$\pm$0.11
Cosine Similarity (Park et al., 2020) -2097.40 -47.26$\pm$2.27 0.52$\pm$0.10
Embedding Weighting (Onan, 2021) -847.56 -29.28$\pm$2.32 0.86$\pm$0.14
Embedding Pruning (Manrique et al., 2023) -846.20 -30.39$\pm$2.76 0.86$\pm$0.04
Ensemble Similarity (Yu et al., 2024) -812.23 -28.64$\pm$3.28 0.89$\pm$0.12
IBIS (proposed) -719.07 -23.17$\pm$3.29 0.93$\pm$0.04
@@ 8 Discussion
Many applications of LLMs are interested in tailoring use cases to individuals, even when little information is known about that individual. While many approaches of individualization exist but have typically relied on advanced machine learning techniques. The method proposed in this work is relatively simple from a mathematical perspective, though there is a strength in its reliance on theories of cognition that underlie human learning and decision making. The result is a simple to understand and easy to implement method of calculating similarities of unseen documents using a cognitive model, which can augment datasets that contain only a small number of decisions.
The general method described here, of augmenting subjective similarity metrics with predicted decisions from a cognitive model, could be applied onto various other scenarios. This includes settings that leverage representations formed of visual information such as $\beta$-Variational Autoencoders (Higgins et al., 2016), which have been related to biological representation formation (Higgins et al., 2021). Overall, we believe that this method is useful for any application where the experience of end-users impacts future decisions.
For instance, in visual learning settings VAEs have been integrated with cognitive models to predict human utility learning of abstract visual information (Malloy and Sims, 2024). Other integrations of Generative AI into cognitive models includes use of LLMs as a knowledge repositories within cognitive models (Kirk et al., 2023). In particular, ConceptNet (Speer et al., 2017) has previously been integrated into a cognitive model for question answering (Huet et al., 2021). Future research should investigate how additional uses of LLMs in integrations of cognitive models can aid in educational settings.
@@ 9 Limitations
The semantic similarity metric suffered from the sparsity of semantic categories in ham emails, additional annotations could raise the performance of this metric and can be explored in future work. However, this ensemble method was partially responsible for the high KDE score of the ensemble method, as it allowed for an integration of both semantic information and embedding similarity. Our IBIS method still outperformed the ensemble method suggesting that this ensemble alone does not address the issues of alternative methods.
One limitation inherent in IBL cognitive models is the time requirements to compare the current instance to all instances in memory. This may make the proposed model unsuitable for applications that rely on large datasets of individual behavior. However, methods in instance compression exist for IBL models (Nguyen et al., 2023). In this setting, we were able to predict individual participant’s decisions fast enough that this was unnecessary.
The specific application we investigated is somewhat unique in that it is based on training human participants to make categorization judgements of textual information of one of two categories. Additionally, the task of annotating whether an email is phishing or ham relies heavily on a small number of features within the email. Namely, if an email contains a link that redirects to a nefarious website, or requests personal information, then it should be labelled as phishing. While students rely on many queues to make their judgements, the annotation is in reality simple. Future work in the area of learning subjective similarity metrics should expand into more complex domains.
10 Acknowledgement
This research was sponsored by the Army Research Office and accomplished under Australia-US MURI Grant Number W911NF-20-S-000, and the AI Research Institutes Program funded by the National Science Foundation under AI Institute for Societal Decision Making (AI-SDM), Award No. 2229881. Compute resources and GPT model credits were provided by the Microsoft Accelerate Foundation Models Research Program grant “Personalized Education with Foundation Models via Cognitive Modeling".